OpenStudy (anonymous):

1. Multiply and simplify. a) (b + 2)(b + 2) b) (x + 3)(x – 3)

OpenStudy (anonymous):

b^2+2b+2b+4 =b^2 + 4b +4 x^2+3x -3x-9 =x^2 - 9

OpenStudy (timaashorty):

\[ a. (b+2)^{2}\] \[b. x ^{2} - 9\]